OpenStudy (anonymous):

Use the elimination method to solve the system. 5x + 8y = –29 7x – 2y = –67

OpenStudy (uri):

5x + 8y = –29 7x – 2y = –67 (multiply by 4) 5x + 8y = –29 28x-8y=-268 ========= 33x=-297 x=-297/33 x=-9

OpenStudy (uri):

7x – 2y = –67 Put the value of x in the original 2 equation 7(-9)-2y=-67 -63-2y=-67 -2y=-67+63 -2y=-4 y=-4/-2 y=4/2 y=2 So x=-9 and y =2 answer.